OpenStudy (anonymous):

Points B, D, and F are midpoints of the sides of triangle ACE. EC=34, AF=9, and DF=19. Find the perimeter of triangle ACE AND TRIANGLE BDF

OpenStudy (anonymous):

AF = 9, AE = 18 EC = 34 DF/AC = EF/EA = 1/2 so AC = 2*DF = 38, so area is 305.8676 area of BDF is area of ACE divided by 4 = 76.4669